Solve the following equations.  17 + 6p = 9 
Solve the following equations.  17 + 6p = 9 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

17+6p=9
6p=9–17
6p=-8
p=-8/6
p=-4/3

Thanks